# Can Finetuing LLMs on Small Human Samples Increase Heterogeneity, Alignment, and Belief-Action Coherence?

### v2

@@ 1 Introduction
The aim of this study is not to enter the discourse of whether off-the-shelf or pre-finetuned LLMs can fully substitute for human respondents, but rather to examine whether finetuning on a small subset of human data, such as that obtainable from a pilot study, can mitigate the limitations documented in prior work. For reasons detailed below, we base our simulations on the behavioral experiment of Hunt, which investigates how information disclosure shapes attackers’ judgments and decisions in a security scenario. Our findings show that LLMs finetuned on small fractions of the human dataset can substantially improve heterogeneity, distributional alignment, and value–action coherence relative to the base LLM. However, despite these improvements, none of the finetuned models reliably reproduce the original regression coefficients or underlying hypothesis tests, indicating that replacing human participants with LLMs remains risky and may introduce non-negligible bias into conclusions. We posit that the most promising use of LLMs in this context is in-silico prototyping for experimental design, helping researchers explore which treatments are likely to yield detectable effects, refine survey instruments, and estimate sample sizes, rather than serving as a direct replacement for human subjects (anthisposition; dillion2023can; binz2025foundation). In sum, we contribute to the literature by showing that fine-tuning on a small sample of human data (30 observations) can substantially mitigate several issues identified in prior work, including the lack of heterogeneity, distributional misalignment, and the value–action gap. Both the original and simulated data will be released for future research upon acceptance.
@@ 4 Methods
Our goal is to explore how well various LLM-based approaches can predict various results derived from this experimental dataset, and in particular how an approach fine-tuned on a subset of this data (e.g. that would come from a pilot study) can improve these predictions. We consider three different approaches, described below. All simulations were conducted using GPT-4.1 (achiam2023gpt) with the default temperature setting of 1.0. The finetuned models use the same base architecture and temperature configuration to ensure comparability.
Finetuning on experimental data
We used two sampling strategies and three dataset sizes, resulting in six models in total. The first strategy uses random sampling, and the second strategy, which we refer to as balanced sampling, draws equal proportions  [^2] We use a round-robin sampling strategy in which we randomly draw one sample from each ethnicity group in sequence until reaching 25% of the dataset. For ethnicity categories with smaller sample sizes (e.g., Other), we include all available observations and then continue the round-robin procedure with the remaining groups to reach the target proportion. of responses from each ethnic subgroup in order to address the systematic misalignment for minority groups noted in prior research. The three dataset sizes correspond to the number of treatment groups included in finetuning. This design allows us to test whether finetuning on only a subset of treatment groups enables the model to generalize to the remaining groups. Specifically, we sample 25 percent of the data from the first treatment group (1a1b), from the first four treatment groups, and from all eight treatment groups. For the multi group settings, we draw 25 percent from each included group. We name each model using its sampling strategy and dataset size. For example, the model finetuned on 25 percent of the first treatment group using random sampling is labeled Random (1a1b), with parallel naming for the remaining random and balanced models.
As outcomes, we examine how finetuning affects heterogeneity, distributional alignment, and the belief–action gap, as well as the accuracy of regression coefficient estimates relative to the regression analysis performed on the human data.
@@ 5.1 Heterogeneity
We find that even a small number of finetuned samples can drastically improve diversity and mitigate homogeneity issues relative to the unfinetuned baseline. Figure 1 shows the distribution of beliefs regarding the likelihood that the defender’s security measures were deployed at Target 1 and both Targets. The analysis of Target 2’s beliefs and the results from all models are presented in Appendix C, and support findings discussed here. In analyzing Target 1 (left panel) in Figure 1, the unfinetuned baseline exhibits a highly concentrated distribution, with most responses clustered in the 30–40 % and 40–50 % bins. In contrast, human responses follow a more bell-shaped pattern, with probability mass spread across multiple bins. This supports prior findings that base LLMs lack response heterogeneity. In an attempt to mitigate this issue, we finetune the model on a small fraction of the human data to examine whether heterogeneity would increase. Specifically, we finetune the model with just 25 % of the data from a single treatment group (the 1a1b treatment group; see Appendix B), equating to 30 of the 929 observations. When we finetune the model on 25 % of the first treatment group, the distribution becomes substantially more dispersed, showing probability mass across all bins and modest peaks around 50–70 %, closely resembling human patterns.
@@ 5.2 Alignment
Given that finetuning on data subsets increases response heterogeneity, we next examine whether this added variation reflects meaningful alignment with human data or merely introduces random noise. To quantify the discrepancy between LLM-simulated and human belief distributions, we use the Jensen–Shannon (JS) distance (lin2002divergence), a symmetric and bounded measure of distributional divergence that has also been applied by Gao to compare LLM and human response distributions. Our results show that all finetuned models substantially reduce the Jensen–Shannon (JS) distance between LLM-simulated data and human responses. Figure 2 reports the JS distance by ethnicity and model. The ethnicity categories correspond to the default classifications used by Prolific, the participant recruitment platform through which the original data were collected. As expected, the unfinetuned baseline exhibits the largest JS distances, consistent with the patterns observed in the previous section. Across all finetuned variants, the JS distance decreases by at least half, indicating improved alignment with human response distributions.
However, the impact of balanced sampling is mixed. Surprisingly, incorporating more demographically diverse data does not necessarily reduce JS distance for minority groups, even though such sampling increases representation. Notably, the lowest JS distance is consistently observed in the White participant category, even for the balanced model, which aligns with prior findings that LLMs most closely reflect the WEIRD population. This pattern emphasizes the persistence of demographic bias in model alignment and highlights the need for more sophisticated sampling or finetuning strategies to achieve equitable representational fidelity.
@@ 5.3 Value Action Gap
While shen2025mind’s approach is theoretically grounded and conceptually novel, the authors do not benchmark their findings against the human value–action gap, which is essential for assessing how closely LLMs approximate human decision-making. In our experimental setting, we define a value–action gap as occurring when a participant, human or LLM, assigns the highest belief probability to a particular target being secured yet still chooses to attack that same target. In reality, if an attacker believes a target it highly secure, it’s unlikely they would attack that target as opposed to selecting a less secure target or selecting not to attack. The unfinetuned model exhibits the largest value–action gap, almost always choosing to attack the target it assigns the highest probability of being secured. In contrast, all finetuned models display substantially smaller gaps. As shown in Figure 3, the value–action gap for the unfinetuned baseline is close to 100%, whereas even models finetuned on only 25% of a single treatment group reduce the gap by more than half. Unlike the alignment results in the previous section, where the relationship between the number of included treatment groups and JS distance was ambiguous, the value–action gap consistently shrinks as more treatment groups are incorporated into finetuning. Models trained on 25% of all eight treatment groups attain discrepancy rates that closely match human participants (23.8% and 24.4% vs. 24.7%).
@@ 5.4 Coefficient Estimation
Given that finetuned models can increase heterogeneity, improve alignment with human data, and reduce the value–action gap, the central question becomes whether these simulated responses can recover the same hypothesis tests and coefficient estimates as the original human experiment. To assess this, we replicate the six regression models reported in the original study. These models fall into three conceptual categories: the first two examine whether beliefs predict an attacker’s decision to abort, the next two predict which target attackers choose, and the final two assess target choice using the normative expected value difference.
@@ 6 Conclusion
Despite successfully mitigating several undesirable properties of the base model, our findings show that none of the finetuned models accurately replicate the regression results from the original study. This limitation indicates that full replacement of human participants remains infeasible and may introduce significant threats to the validity of empirical conclusions. Instead, we argue that LLM-based simulation is most promising as a tool for in-silico prototyping, helping researchers anticipate which treatments may produce detectable effects, refine survey instruments, and conduct power or sample size analyses, rather than as a direct substitute for human subjects (anthisposition; dillion2023can; binz2025foundation).
@@ 7 Discussion
Several mechanisms may explain the misalignments between LLM-simulated and human responses. wang2025large argue that LLMs often engage in misportraying, rather than representing the authentic behaviors of a given demographic group, models reproduce what out-group members believe those in-group members would do. Since demographic identity is rarely encoded directly in the text on which LLMs are trained, explicit mentions of identity may instead express stereotypes by out-group authors. liu2024large offer a related explanation, and it aligns with the sociolingustic theoretical model introduced by [https://arxiv.org/abs/2502.12057]. They study LLM performance on datasets involving risky decision-making and find that LLMs align more closely with people’s expectations of how others behave than with how individuals actually behave. Since people tend to assume that others are more rational than themselves, LLMs may encode these perceptions rather than genuine human behavior. These findings suggest that LLM decisions may be shaped less by representations of actual in-group behavior and more by culturally shared beliefs about what “typical” behavior should look like. Additional factors may also contribute to misalignment. The training data themselves may contain survey responses affected by social desirability bias (crowne1960new), meaning that models learn socially acceptable patterns of responding rather than authentic responses. More directly, reinforcement learning from human feedback (RLHF) (ouyang2022training) optimizes models to behave in ways that humans prefer LLMs to behave (dahlgren2025helpful), not necessarily in ways that mirror empirical human data.

### v3

@@ 1 Introduction
The aim of this study is not to enter the discourse of whether off-the-shelf or pre-finetuned LLMs can fully substitute for human respondents, but rather to examine whether finetuning on a small subset of human data, such as that obtainable from a pilot study, can mitigate the limitations documented in prior work. For reasons detailed below, we base our simulations on the behavioral experiment of Hunt, which investigates how information disclosure shapes attackers’ judgments and decisions in a security scenario. Our findings show that LLMs finetuned on small fractions of the human dataset can substantially improve heterogeneity, distributional alignment, and value–action coherence relative to the base LLM. However, despite these improvements, none of the finetuned models reliably reproduce the original regression coefficients or underlying hypothesis tests, indicating that replacing human participants with LLMs remains risky and may introduce non-negligible bias into conclusions. Hence, we caution against the naive use of base models for experimental design and simulation, even when augmented with sociodemographic prompting, they tend to produce overly tight variance. Such compression can yield overconfident power analyses and lead to underpowered studies. In contrast, we show that finetuning on as few as 30 human observations substantially improves distributional alignment, producing variation that more closely resembles human data and yielding more realistic simulated responses. This makes finetuned models better suited for certain stages of survey design and pretesting than base LLMs. In sum, we contribute to the literature by showing that finetuning on a small sample of human data (30 observations) can substantially mitigate several issues identified in prior work, including the lack of heterogeneity, distributional misalignment, and the value–action gap. Both the original and simulated data will be released for future research upon acceptance.
@@ 4 Methods
Our goal is to explore how well various LLM-based approaches can predict various results derived from this experimental dataset, and in particular how an approach finetuned on a subset of this data (e.g. that would come from a pilot study) can improve these predictions. We consider three different approaches, described below. All simulations were conducted using GPT-4.1 (achiam2023gpt) with the default temperature setting of 1.0. The finetuned models use the same base architecture and temperature configuration to ensure comparability. We mimic most prior work discussed above and first ran simulations using an unfinetuned GPT model. We incorporated participant demographics as role prompts to approximate more realistic responses. For finetuning, we used two sampling strategies and three dataset sizes, resulting in six models in total. The first strategy uses random sampling, and the second strategy, which we refer to as balanced sampling, draws equal proportions  [^2] We use a round-robin sampling strategy in which we randomly draw one sample from each ethnicity group in sequence until reaching 25% of the dataset. For ethnicity categories with smaller sample sizes (e.g., Other), we include all available observations and then continue the round-robin procedure with the remaining groups to reach the target proportion. of responses from each ethnic subgroup in order to address the systematic misalignment for minority groups noted in prior research. The three dataset sizes correspond to the number of treatment groups included in finetuning. This design allows us to test whether finetuning on only a subset of treatment groups enables the model to generalize to the remaining groups. Specifically, we sample 25% of the data from the first treatment group (1a1b), from the first four treatment groups, and from all eight treatment groups. For the multi group settings, we draw 25% from each included group. We name each model using its sampling strategy and dataset size. For example, the model finetuned on 25% of the first treatment group using random sampling is labeled Random (1a1b), with parallel naming for the remaining random and balanced models.
As outcomes, we examine how finetuning affects heterogeneity with just 25 % of the data from a single treatment group (the 1a1b treatment group; see Appendix B). To quantify distributional alignment, we use the Jensen–Shannon (JS) distance (lin2002divergence), a symmetric and bounded measure of distributional divergence that has also been applied by Gao to compare LLM and human response distributions. For belief–action gap, we measure the conditional proportion of LLMs choosing to attack the target that they think has the highest probability that security is in place and compare it to humans. Lastly, we compute the accuracy of regression coefficient estimates relative to the regression analysis performed on the human data.
@@ 5.1 Heterogeneity
We find that even a small number of finetuned samples can drastically improve diversity and mitigate homogeneity issues relative to the unfinetuned baseline. Figure 1 shows the distribution of beliefs regarding the likelihood that the defender’s security measures were deployed at Target 1 and both Targets. The analysis of Target 2’s beliefs and the results from all models are presented in Appendix C, and support findings discussed here. In analyzing Target 1 (left panel) in Figure 1, the unfinetuned baseline exhibits a highly concentrated distribution, with most responses clustered in the 30–40 % and 40–50 % bins. In contrast, human responses follow a more bell-shaped pattern, with probability mass spread across multiple bins. This supports prior findings that base LLMs lack response heterogeneity. However, when we finetune the model on 25 % of the first treatment group using the random sampling strategy, the distribution becomes substantially more dispersed, showing probability mass across all bins and modest peaks around 50–70 %, closely resembling human patterns.
@@ 5.2 Alignment
Given that finetuning on data subsets increases response heterogeneity, we next examine whether this added variation reflects meaningful alignment with human data or merely introduces random noise. Our results show that all finetuned models substantially reduce the Jensen–Shannon (JS) distance between LLM-simulated data and human responses. Figure 2 reports the JS distance by ethnicity and model. The ethnicity categories correspond to the default classifications used by Prolific, the participant recruitment platform through which the original data were collected. As expected, the unfinetuned baseline exhibits the largest JS distances, consistent with the patterns observed in the previous section. Across all finetuned variants, the JS distance decreases by at least half, indicating improved alignment with human response distributions.
However, the impact of balanced sampling is mixed. Surprisingly, incorporating more demographically diverse data does not necessarily reduce JS distance for minority groups, even though such sampling increases representation. Notably, the lowest JS distance is consistently observed in the White participant category, even for the balanced model, which aligns with prior findings that LLMs most closely reflect the WEIRD population. This pattern emphasizes the persistence of demographic bias in model alignment and highlights the need for more sophisticated sampling or finetuning strategies to achieve equitable representational fidelity. Indeed, our ablation study (see Appendix E) shows that adding demographic information does not meaningfully increase alignment in our context.
@@ 5.3 Value Action Gap
While the original value-action gap paper (shen2025mind)’s approach is theoretically grounded and conceptually novel, the authors do not benchmark their findings against the human value–action gap, which is essential for assessing how closely LLMs approximate human decision-making. In our experimental setting, we define a value–action gap as occurring when a participant, human or LLM, assigns the highest belief probability to a particular target being secured yet still chooses to attack that same target. In reality, if an attacker believes a target it highly secure, it’s unlikely they would attack that target as opposed to selecting a less secure target or selecting not to attack. The unfinetuned model exhibits the largest value–action gap, almost always choosing to attack the target it assigns the highest probability of being secured. In contrast, all finetuned models display substantially smaller gaps. As shown in Figure 3, the value–action gap for the unfinetuned baseline is close to 100%, whereas even models finetuned on only 25% of a single treatment group reduce the gap by more than half. Unlike the alignment results in the previous section, where the relationship between the number of included treatment groups and JS distance was ambiguous, the value–action gap consistently shrinks as more treatment groups are incorporated into finetuning. Models trained on 25% of all eight treatment groups attain discrepancy rates that closely match human participants (23.8% and 24.4% vs. 24.7%).
@@ 5.4 Coefficient Estimation
Given that finetuned models can increase heterogeneity, improve alignment with human data, and reduce the value–action gap, the natural next step becomes whether these simulated responses can recover the same hypothesis tests and coefficient estimates as the original human experiment. To assess this, we replicate the six regression models reported in the original study. These models fall into three conceptual categories: the first two examine whether beliefs predict an attacker’s decision to abort, the next two predict which target attackers choose, and the final two assess target choice using the normative expected value difference.
@@ 6 Conclusion
Despite successfully mitigating several undesirable properties of the base model, our findings show that none of the finetuned models accurately replicate the regression results from the original study. This limitation indicates that full replacement of human participants remains infeasible and may introduce significant threats to the validity of empirical conclusions. Instead, we argue that LLMs finetuned on small human samples, as few as 30 observations, are better suited for in silico prototyping than base LLMs. Such models can assist researchers in refining survey instruments and conducting power or sample size analyses, but should be viewed as complements to, rather than substitutes for, human subjects (anthisposition; dillion2023can; binz2025foundation).
@@ 7 Discussion
Several mechanisms may explain the misalignments between LLM-simulated and human responses. wang2025large argue that LLMs often engage in misportraying, rather than representing the authentic behaviors of a given demographic group, models reproduce what out-group members believe those in-group members would do. Since demographic identity is rarely encoded directly in the text on which LLMs are trained, explicit mentions of identity may instead express stereotypes by out-group authors. liu2024large offer a related explanation, and it aligns with the sociolingustic theoretical model introduced by zhou2025culture. They study LLM performance on datasets involving risky decision-making and find that LLMs align more closely with people’s expectations of how others behave than with how individuals actually behave. Since people tend to assume that others are more rational than themselves, LLMs may encode these perceptions rather than genuine human behavior. These findings suggest that LLM decisions may be shaped less by representations of actual in-group behavior and more by culturally shared beliefs about what “typical” behavior should look like. Additional factors may also contribute to misalignment. The training data themselves may contain survey responses affected by social desirability bias (crowne1960new), meaning that models learn socially acceptable patterns of responding rather than authentic responses. More directly, reinforcement learning from human feedback (RLHF) (ouyang2022training) optimizes models to behave in ways that humans prefer LLMs to behave (dahlgren2025helpful), not necessarily in ways that mirror empirical human data.
@@ Appendix D Group-Level JS
Appendix E Ablation JS
Figure 7 presents an ablation study conducted on one of our finetuned models. Random – No Demographic refers to a variant trained without participants’ background information. Random – No Decision and Random – No Belief refer to variants that exclude participants’ reported final decisions and belief responses, respectively, during finetuning. Due to budget constraints, we perform this ablation on a single representative model (Random (4 Groups)) rather than across all finetuned models in the manuscript. As shown in Figure 7, the Jensen–Shannon (JS) distances for Random – No Demographic are similar to those of the original finetuned model across ethnic subgroups, suggesting that including demographic background information alone does not meaningfully reduce subgroup misalignment in belief distributions. In contrast, Random – No Belief exhibits larger JS distances than the other finetuned variants, indicating increased misalignment when decision information is excluded, nevertheless, its performance remains substantially better than that of the unfinetuned baseline. Finally, Random – No Decision performs comparably to the original model, implying that adding decision information does not indirectly improve belief alignment when belief data themselves are omitted during finetuning.
Figure 7: Ablation study of a finetuned LLM showing Jensen–Shannon (JS) distances between simulated and human belief distributions by ethnicity (lower values indicate closer alignment). “Random – No Demographic,” “Random – No Decision,” and “Random – No Belief” exclude demographic background, decision responses, and belief responses, respectively, during finetuning. The unfinetuned baseline is shown for comparison.